\begin{thebibliography}{10}

\bibitem{APPLEBY}
D.~M. Appleby.
\newblock Symmetric informationally complete-positive operator valued measures
  and the extended {C}lifford group.
\newblock {\em J. Math. Phys.}, 46(5):052107, 29, 2005.

\bibitem{APPLEBYBENGTSSONFLAMMIAGOYENECHE}
Marcus Appleby, Ingemar Bengtsson, Steven Flammia, and Dardo Goyeneche.
\newblock Tight frames, {H}adamard matrices and {Z}auner's conjecture.
\newblock {\em J. Phys. A}, 52(29):295301, 26, 2019.

\bibitem{APPLEBYFLAMMIAMCCONNELLYARD}
Marcus Appleby, Steven Flammia, Gary McConnell, and Jon Yard.
\newblock S{IC}s and algebraic number theory.
\newblock {\em Found. Phys.}, 47(8):1042--1059, 2017.

\bibitem{BALLADRAXLERKEEVASHSUDAKOV2}
Igor Balla, Felix Draxler, Peter Keevash, and Benny Sudakov.
\newblock Equiangular lines and subspaces in {E}uclidean spaces.
\newblock {\em Electronic Notes in Discrete Mathematics}, 61:85--91, 2017.

\bibitem{BALLADRAXLERKEEVASHSUDAKOV}
Igor Balla, Felix Dr\"{a}xler, Peter Keevash, and Benny Sudakov.
\newblock Equiangular lines and spherical codes in {E}uclidean space.
\newblock {\em Invent. Math.}, 211(1):179--212, 2018.

\bibitem{BALLASUDAKOV}
Igor Balla and Benny Sudakov.
\newblock Equiangular subspaces in {E}uclidean spaces.
\newblock {\em Discrete Comput. Geom.}, 61(1):81--90, 2019.

\bibitem{BARGYU}
Alexander Barg and Wei-Hsuan Yu.
\newblock New bounds for equiangular lines.
\newblock In {\em Discrete geometry and algebraic combinatorics}, volume 625 of
  {\em Contemp. Math.}, pages 111--121. Amer. Math. Soc., Providence, RI, 2014.

\bibitem{BROUWERHAEMERS}
Andries~E. Brouwer and Willem~H. Haemers.
\newblock {\em Spectra of graphs}.
\newblock Universitext. Springer, New York, 2012.

\bibitem{BUKH}
Boris Bukh.
\newblock Bounds on equiangular lines and on related spherical codes.
\newblock {\em SIAM J. Discrete Math.}, 30(1):549--554, 2016.

\bibitem{CONWAYHARDINSLOANE}
John~H. Conway, Ronald~H. Hardin, and Neil J.~A. Sloane.
\newblock Packing lines, planes, etc.: packings in {G}rassmannian spaces.
\newblock {\em Experiment. Math.}, 5(2):139--159, 1996.

\bibitem{COUTINHOGODSILSHIRAZIZHAN}
G.~Coutinho, C.~Godsil, H.~Shirazi, and H.~Zhan.
\newblock Equiangular lines and covers of the complete graph.
\newblock {\em Linear Algebra Appl.}, 488:264--283, 2016.

\bibitem{DECAEN}
D.~de~Caen.
\newblock Large equiangular sets of lines in {E}uclidean space.
\newblock {\em Electron. J. Combin.}, 7:Research Paper 55, 3, 2000.

\bibitem{DELAATMACHADODEOLIVEIRAFERNANDAVELLENTIN}
David de~Laat, Fabr\'{\i}cio~Caluza Machado, Fernando~M\'{a}rio
  de~Oliveira~Filho, and Frank Vallentin.
\newblock {$k$}-point semidefinite programming bounds for equiangular lines.
\newblock {\em Math. Program.}, 194(1-2, Ser. A):533--567, 2022.

\bibitem{GIPSON}
Philip~M. Gipson.
\newblock Invariant basis number for {$C^*$}-algebras.
\newblock {\em Illinois J. Math.}, 59(1):85--98, 2015.

\bibitem{GLAZYRINYU}
Alexey Glazyrin and Wei-Hsuan Yu.
\newblock Upper bounds for {$s$}-distance sets and equiangular lines.
\newblock {\em Adv. Math.}, 330:810--833, 2018.

\bibitem{GODSILROY}
Chris Godsil and Aidan Roy.
\newblock Equiangular lines, mutually unbiased bases, and spin models.
\newblock {\em European J. Combin.}, 30(1):246--262, 2009.

\bibitem{GODSILROYLE}
Chris Godsil and Gordon Royle.
\newblock {\em Algebraic graph theory}, volume 207 of {\em Graduate Texts in
  Mathematics}.
\newblock Springer-Verlag, New York, 2001.

\bibitem{GREAVESKOOLENMUNEMASASZOLLOSO}
Gary Greaves, Jacobus~H. Koolen, Akihiro Munemasa, and Ferenc Sz\"{o}ll\H{o}si.
\newblock Equiangular lines in {E}uclidean spaces.
\newblock {\em J. Combin. Theory Ser. A}, 138:208--235, 2016.

\bibitem{GREAVES7}
Gary R.~W. Greaves.
\newblock Equiangular line systems and switching classes containing regular
  graphs.
\newblock {\em Linear Algebra Appl.}, 536:31--51, 2018.

\bibitem{GREAVESIVERSONJASPERMIXON}
Gary R.~W. Greaves, Joseph~W. Iverson, John Jasper, and Dustin~G. Mixon.
\newblock Frames over finite fields: basic theory and equiangular lines in
  unitary geometry.
\newblock {\em Finite Fields Appl.}, 77:Paper No. 101954, 41, 2022.

\bibitem{GREAVESIVERSONJASPERMIXON2}
Gary R.~W. Greaves, Joseph~W. Iverson, John Jasper, and Dustin~G. Mixon.
\newblock Frames over finite fields: equiangular lines in orthogonal geometry.
\newblock {\em Linear Algebra Appl.}, 639:50--80, 2022.

\bibitem{GREAVESSYATRIADI2}
Gary R.~W. Greaves and Jeven Syatriadi.
\newblock Real equiangular lines in dimension 18 and the {J}acobi identity for
  complementary subgraphs.
\newblock {\em J. Combin. Theory Ser. A}, 201:Paper No. 105812, 27, 2024.

\bibitem{GREAVESSYATRIADIYATSYNA2}
Gary R.~W. Greaves, Jeven Syatriadi, and Pavlo Yatsyna.
\newblock Equiangular lines in low dimensional {E}uclidean spaces.
\newblock {\em Combinatorica}, 41(6):839--872, 2021.

\bibitem{GREAVSSYATRIADIYATSYNA}
Gary R.~W. Greaves, Jeven Syatriadi, and Pavlo Yatsyna.
\newblock Equiangular lines in {E}uclidean spaces: dimensions 17 and 18.
\newblock {\em Math. Comp.}, 92(342):1867--1903, 2023.

\bibitem{HAANTJES}
J.~Haantjes.
\newblock Equilateral point-sets in elliptic two- and three-dimensional spaces.
\newblock {\em Nieuw Arch. Wiskunde (2)}, 22:355--362, 1948.

\bibitem{HOFFMANSOLAZZO}
Thomas~R. Hoffman and James~P. Solazzo.
\newblock Complex equiangular tight frames and erasures.
\newblock {\em Linear Algebra Appl.}, 437(2):549--558, 2012.

\bibitem{JEDWABWIEBE}
Jonathan Jedwab and Amy Wiebe.
\newblock Large sets of complex and real equiangular lines.
\newblock {\em J. Combin. Theory Ser. A}, 134:98--102, 2015.

\bibitem{JEDWABWIEBE2}
Jonathan Jedwab and Amy Wiebe.
\newblock Constructions of complex equiangular lines from mutually unbiased
  bases.
\newblock {\em Des. Codes Cryptogr.}, 80(1):73--89, 2016.

\bibitem{JIANGPOLYANSKII}
Zilin Jiang and Alexandr Polyanskii.
\newblock Forbidden subgraphs for graphs of bounded spectral radius, with
  applications to equiangular lines.
\newblock {\em Israel J. Math.}, 236(1):393--421, 2020.

\bibitem{JIANGTIDORYAOZHANGZAO}
Zilin Jiang, Jonathan Tidor, Yuan Yao, Shengtong Zhang, and Yufei Zhao.
\newblock Equiangular lines with a fixed angle.
\newblock {\em Ann. of Math. (2)}, 194(3):729--743, 2021.

\bibitem{KAPLANSKY}
Irving Kaplansky.
\newblock Modules over operator algebras.
\newblock {\em Amer. J. Math.}, 75:839--858, 1953.

\bibitem{KINGTANG}
Emily~J. King and Xiaoxian Tang.
\newblock New upper bounds for equiangular lines by pillar decomposition.
\newblock {\em SIAM J. Discrete Math.}, 33(4):2479--2508, 2019.

\bibitem{KOORNWINDER}
Tom~H. Koornwinder.
\newblock A note on the absolute bound for systems of lines.
\newblock {\em Indag. Math.}, 38(2):152--153, 1976.

\bibitem{KOPP}
Gene~S. Kopp.
\newblock S{IC}-{POVM}s and the {S}tark conjectures.
\newblock {\em Int. Math. Res. Not. IMRN}, (18):13812--13838, 2021.

\bibitem{MAHESH}
K.~Mahesh Krishna.
\newblock $p$-adic equiangular lines and $p$-adic van {L}int–{S}eidel
  relative bound.
\newblock {\em Bulletin Polish Acad. Sci. Math}, 2014.

\bibitem{LEMMENSSEIDEL}
P.~W.~H. Lemmens and J.~J. Seidel.
\newblock Equiangular lines.
\newblock {\em J. Algebra}, 24:494--512, 1973.

\bibitem{LINYU2}
Yen-Chi~Roger Lin and Wei-Hsuan Yu.
\newblock Equiangular lines and the {L}emmens-{S}eidel conjecture.
\newblock {\em Discrete Math.}, 343(2):111667, 18, 2020.

\bibitem{LINYU}
Yen-chi~Roger Lin and Wei-Hsuan Yu.
\newblock Saturated configuration and new large construction of equiangular
  lines.
\newblock {\em Linear Algebra Appl.}, 588:272--281, 2020.

\bibitem{MALOZEMOPEVNYI}
V.~N. Malozemov and A.~B. Pevnyi.
\newblock Equiangular tight frames.
\newblock {\em J. Math. Sci. (N.Y.)}, 157(6):789--815, 2009.

\bibitem{NEUMAIER}
A.~Neumaier.
\newblock Graph representations, two-distance sets, and equiangular lines.
\newblock {\em Linear Algebra Appl.}, 114/115:141--156, 1989.

\bibitem{OKUDAYU}
Takayuki Okuda and Wei-Hsuan Yu.
\newblock A new relative bound for equiangular lines and nonexistence of tight
  spherical designs of harmonic index 4.
\newblock {\em European J. Combin.}, 53:96--103, 2016.

\bibitem{PASCHKE}
William~L. Paschke.
\newblock Inner product modules over {$B^{\ast} $}-algebras.
\newblock {\em Trans. Amer. Math. Soc.}, 182:443--468, 1973.

\bibitem{RENESBLUMEKOHOUTSCOTTCAVES}
Joseph~M. Renes, Robin Blume-Kohout, A.~J. Scott, and Carlton~M. Caves.
\newblock Symmetric informationally complete quantum measurements.
\newblock {\em J. Math. Phys.}, 45(6):2171--2180, 2004.

\bibitem{RIEFFEL}
Marc~A. Rieffel.
\newblock Induced representations of {$C^{\ast} $}-algebras.
\newblock {\em Advances in Math.}, 13:176--257, 1974.

\bibitem{SCOTT}
A.~J. Scott.
\newblock Tight informationally complete quantum measurements.
\newblock {\em J. Phys. A}, 39(43):13507--13530, 2006.

\bibitem{SEIDELHANDBOOK}
J.~J. Seidel.
\newblock Discrete non-{E}uclidean geometry.
\newblock In {\em Handbook of incidence geometry}, pages 843--920.
  North-Holland, Amsterdam, 1995.

\bibitem{STROHMER}
Thomas Strohmer.
\newblock A note on equiangular tight frames.
\newblock {\em Linear Algebra Appl.}, 429(1):326--330, 2008.

\bibitem{SUSTIKTROPPDHILLON}
M\'{a}ty\'{a}s~A. Sustik, Joel~A. Tropp, Inderjit~S. Dhillon, and Robert~W.
  Heath, Jr.
\newblock On the existence of equiangular tight frames.
\newblock {\em Linear Algebra Appl.}, 426(2-3):619--635, 2007.

\bibitem{VANLINTSEIDEL}
J.~H. van Lint and J.~J. Seidel.
\newblock Equilateral point sets in elliptic geometry.
\newblock {\em Indag. Math.}, 28:335--348, 1966.
\newblock Nederl. Akad. Wetensch. Proc. Ser. A {{\bf{6}}9}.

\bibitem{WALDONBOOK}
Shayne F.~D. Waldron.
\newblock {\em An introduction to finite tight frames}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user/Springer, New
  York, 2018.

\bibitem{WEGGEOLSEN}
N.~E. Wegge-Olsen.
\newblock {\em {$K$}-theory and {$C^*$}-algebras: {A} friendly approach}.
\newblock Oxford Science Publications. The Clarendon Press, Oxford University
  Press, New York, 1993.

\bibitem{YU}
Wei-Hsuan Yu.
\newblock New bounds for equiangular lines and spherical two-distance sets.
\newblock {\em SIAM J. Discrete Math.}, 31(2):908--917, 2017.

\bibitem{ZAUNER}
Gerhard Zauner.
\newblock Quantum designs: foundations of a noncommutative design theory.
\newblock {\em Int. J. Quantum Inf.}, 9(1):445--507, 2011.

\bibitem{ZHAO}
Yufei Zhao.
\newblock Equiangular lines and eigenvalue multiplicities.
\newblock {\em Notices Amer. Math. Soc.}, 71(9):1151--1159, 2024.

\end{thebibliography}

\documentclass{amsart}
\usepackage{setspace}
\usepackage{a4}
\usepackage{amsthm}
\usepackage{latexsym}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{cite}
\usepackage{enumerate}
\usepackage{amssymb}
\usepackage{hyperref}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage[mathscr]{euscript}
\usepackage{mathtools}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{definition}[theorem]{Definition}

\newtheorem{problem}[theorem]{Problem}

\title{This is the title}
\usepackage{fancyhdr}

\begin{document}
\hrule\hrule\hrule\hrule\hrule
\vspace{0.3cm}	
\begin{center}
{\bf{Noncommutative Equiangular Lines: van Lint-Seidel Relative and Gerzon Universal Bounds}}\\
\vspace{0.3cm}
\hrule\hrule\hrule\hrule\hrule
\vspace{0.3cm}
\textbf{K. Mahesh Krishna}\\
School of Mathematics and Natural Sciences\\
Chanakya University Global Campus\\
NH-648, Haraluru Village\\
Devanahalli Taluk, 	Bengaluru  Rural District\\
Karnataka State, 562 110, India \\
Email: kmaheshak@gmail.com\\

Date: \today
\end{center}

\hrule\hrule
\vspace{0.5cm}
\textbf{Abstract}: We introduce the notion of noncommutative equiangular lines and derive noncommutative versions of  fundamental van Lint-Seidel relative and Gerzon universal bounds.

\textbf{Keywords}: Equiangular lines, C*-algebra, Hilbert C*-module.

\textbf{Mathematics Subject Classification (2020)}: 46L05, 46L08.\\

\hrule

\hrule
\section{Introduction}
Let $d \in \mathbb{N}$ and $\gamma \in  [0,1]$. Recall that  a collection $\{\tau_j\}_{j=1}^n$  of unit vectors  in $\mathbb{R}^d$ or $\mathbb{C}^d$	is said to be \textbf{$\gamma$-equiangular  lines} \cite{LEMMENSSEIDEL, HAANTJES}  if 
\begin{align*}
	|\langle \tau_j, \tau_k\rangle|=\gamma, \quad \forall 1\leq j, k \leq n, j \neq k.
\end{align*}	
Two fundamental problems associated with equiangular lines are the following.
\begin{problem}\label{1} 
	Given $d \in \mathbb{N}$ and $\gamma \in  [0,1]$, what is the upper bound on $n$ such that there exists a collection $\{\tau_j\}_{j=1}^n$  of  $\gamma$-equiangular lines in $\mathbb{R}^d$?
\end{problem}
\begin{problem}\label{2} 
	For which parameters $d, n \in \mathbb{N}$ and $\gamma \in  [0,1]$, there exists a collection $\{\tau_j\}_{j=1}^n$  of  $\gamma$-equiangular lines in $\mathbb{R}^d$? If exists, how to construct them?
\end{problem}
Two  answers to Problem (\ref{1}) which are   fundamental driving force in the study of equiangular lines are the following results of van Lint and Seidel \cite{VANLINTSEIDEL, LEMMENSSEIDEL} and Gerzon \cite{WALDONBOOK}.
\begin{theorem} \cite{VANLINTSEIDEL, LEMMENSSEIDEL, GLAZYRINYU, GODSILROYLE} \label{VS} (\textbf{van Lint-Seidel Relative Bound}) 
	Let $\{\tau_j\}_{j=1}^n$  	be  $\gamma$-equiangular  lines in $\mathbb{R}^d$ or $\mathbb{C}^d$. 
	Then 
	\begin{align*}
		n(1-d\gamma^2) \leq d(1-\gamma^2).
	\end{align*}
	In particular, if 
	\begin{align*}
		\gamma < \frac{1}{\sqrt{d}},
	\end{align*}
	then 
	\begin{align}\label{LS}
		n\leq \frac{d(1-\gamma^2)}{1-d\gamma^2}.
	\end{align}
\end{theorem}
\begin{theorem} \cite{WALDONBOOK} \label{GBT}
	(\textbf{Gerzon Universal Bound}) 
	\begin{enumerate}[\upshape(i)]
		\item 
		Let $\{\tau_j\}_{j=1}^n$  	be  $\gamma$-equiangular  lines in $\mathbb{R}^d$. 
		Then 
		\begin{align}\label{GB}
			n \leq \frac{d(d+1)}{2}.
		\end{align}	
		\item 
		Let $\{\tau_j\}_{j=1}^n$  	be  $\gamma$-equiangular  lines in $\mathbb{C}^d$. 
		Then 
		\begin{align}\label{CGB}
			n \leq d^2.
		\end{align}	
	\end{enumerate}
\end{theorem}
Bounds in Inequalities  (\ref{LS}),  (\ref{GB}) and (\ref{CGB}) are  improved for various special values of $\gamma$, $n$ and $d$ (including asymptotic). Regarding Problem (\ref{2}), various special cases are solved but full generality remains open. We refer \cite{BUKH, JIANGTIDORYAOZHANGZAO, BALLADRAXLERKEEVASHSUDAKOV, BARGYU, DECAEN, MALOZEMOPEVNYI, NEUMAIER, JIANGPOLYANSKII, DELAATMACHADODEOLIVEIRAFERNANDAVELLENTIN, GREAVESKOOLENMUNEMASASZOLLOSO, GREAVSSYATRIADIYATSYNA, OKUDAYU, YU, KINGTANG, CONWAYHARDINSLOANE, GREAVESSYATRIADI2, GREAVESSYATRIADIYATSYNA2, SEIDELHANDBOOK, GREAVES7, KOORNWINDER, LINYU, LINYU2, ZHAO, COUTINHOGODSILSHIRAZIZHAN}
 for a look on these achievements in the real case and \cite{JEDWABWIEBE, JEDWABWIEBE2, GODSILROY, SUSTIKTROPPDHILLON, STROHMER, HOFFMANSOLAZZO, BROUWERHAEMERS}
    in the complex case. In the real case, it is known that equality is not always attained for  every $d$ in Inequality (\ref{GB}).   Celebrated Zauner conjecture says that  there exists a   $d^2$ $\frac{1}{\sqrt{d+1}}$-equiangular lines in $\mathbb{C}^d$ for every $d$ satisfying equality in Inequality (\ref{CGB}) \cite{ZAUNER, RENESBLUMEKOHOUTSCOTTCAVES, APPLEBY, APPLEBYBENGTSSONFLAMMIAGOYENECHE, KOPP, APPLEBYFLAMMIAMCCONNELLYARD, SCOTT}.

We wish to record that there is  a vector space version of equiangular lines \cite{GREAVESIVERSONJASPERMIXON2, GREAVESIVERSONJASPERMIXON}  and  there is also a notion of  equiangular subspaces of Euclidean spaces \cite{BALLASUDAKOV, BALLADRAXLERKEEVASHSUDAKOV2}. Recently, p-adic version of equiangular lines have been introduced \cite{MAHESH}.

As noncommutative geometry plays major role in modern Mathematics, our fundamental objective is to initiate a study of noncommutative equiangular lines. As most important results in the theory of equiangular lines are van Lint-Seidel relative and Gerzon universal bounds, we derive noncommutative versions of them. 

\section{Noncommutative Equiangular Lines}
We want the notion of Hilbert C*-modules. They are first introduced by Kaplansky \cite{KAPLANSKY} for modules over commutative C*-algebras and later developed for modules over arbitrary C*-algebras by Paschke  \cite{PASCHKE} and Rieffel \cite{RIEFFEL}. A friendly introduction is given in the book \cite{WEGGEOLSEN}. We want only standard Hilbert C*-module. Let $\mathcal{A}$ be a  unital C*-algebra with identity $1$.
 For $d\in \mathbb{N}$, let $\mathcal{A}^d$ be the standard left Hilbert C*-module over $\mathcal{A}$ with inner product 
\begin{align*}
	\langle (a_j)_{j=1}^d, (b_j)_{j=1}^d\rangle \coloneqq \sum_{j=1}^{d}a_jb_j^*, \quad \forall (a_j)_{j=1}^d, (b_j)_{j=1}^d \in \mathcal{A}^d.
\end{align*}
Hence norm on $\mathcal{A}^d$ is 
\begin{align*}
	\|(a_j)_{j=1}^d\|_2\coloneqq \left\|\sum_{j=1}^{d}a_ja_j^*\right\|^\frac{1}{2}, \quad \forall (a_j)_{j=1}^d \in \mathcal{A}^d.
\end{align*}
 We  introduce two notions of equiangular lines for modules.
\begin{definition}\label{A}
Let $a \in \mathcal{A}$ be a positive	element.   A collection $\{\tau_j\}_{j=1}^n$ in  $\mathcal{A}^d$  is said to be \textbf{modular $a$-equiangular  lines} if following conditions hold.
\begin{enumerate}[\upshape(i)]
	\item $	\langle \tau_j, \tau_j \rangle =1,  \forall 1\leq j \leq n$.
	\item $	\langle\tau_j, \tau_k \rangle\langle\tau_k, \tau_j \rangle=a,  \forall 	1\leq j , k \leq n , j\neq k.$
\end{enumerate}
\end{definition}
\begin{definition}
	Let $\gamma\in [0,1].$   A collection $\{\tau_j\}_{j=1}^n$ in  $\mathcal{A}^d$  is said to be \textbf{modular $\gamma$-norm-equiangular  lines} if following conditions hold.
	\begin{enumerate}[\upshape(i)]
		\item $	\langle \tau_j, \tau_j \rangle =1,  \forall 1\leq j \leq n$.
		\item $	\|\langle\tau_j, \tau_k \rangle\|=\gamma,  \forall 	1\leq j , k \leq n , j\neq k.$
	\end{enumerate}
\end{definition}
It is clear that if $\{\tau_j\}_{j=1}^n$ is modular  $a$-equiangular  lines, then $\{\tau_j\}_{j=1}^n$ is modular  $\sqrt{\|a\|}$-norm-equiangular  lines. We now derive two modular versions  of Theorem \ref{VS} for modules over commutative C*-algebras.
\begin{theorem} (\textbf{Modular van Lint-Seidel Relative Bound}) \label{MVS}
Let $\mathcal{A}$ be a commutative unital C*-algebra. 	Let $\{\tau_j\}_{j=1}^n$  	be modular  $a$-equiangular  lines in $\mathcal{A}^d$.
	Then 
	\begin{align*}
		n(1-da) \leq d(1-a).
	\end{align*}
	In particular, if $1-da$ is invertible, then 
	\begin{align*}
		n\leq \frac{d(1-a)}{1-da}\leq d \left\|\frac{1-a}{1-da}\right\|.
	\end{align*}
	\end{theorem}
\begin{proof}
Let $\tau_j\coloneqq(a_1^{(j)}, a_2^{(j)}, \dots, a_d^{(j)}) $ for all $1\leq j\leq n$. Define 
	\begin{align*}
		B\coloneqq \sum_{j=1}^n\sum_{k=1}^n\langle \tau_j, \tau_k\rangle \langle \tau_k, \tau_j\rangle =n+(n^2-n)a.
	\end{align*}
Then using Cauchy-Schwarz inequality in Hilbert C*-modules, 
\begin{align*}
	B&= \sum_{j=1}^n\sum_{k=1}^n\left(\sum_{r=1}^{d}a_r^{(j)}(a_r^{(k)})^*\right)	\left(\sum_{s=1}^{d}a_s^{(k)}(a_s^{(j)})^*\right)\\
	&=\sum_{r, s=1}^{d}\left(\sum_{j=1}^na_r^{(j)}(a_s^{(j)})^*\right)\left(\sum_{k=1}^na_r^{(k)}(a_s^{(k)})^*\right)^*\\
	&=\sum_{r=1}^{d}\left(\sum_{j=1}^na_r^{(j)}(a_r^{(j)})^*\right)\left(\sum_{k=1}^na_r^{(k)}(a_r^{(k)})^*\right)^*+\sum_{r, s=1, r\neq s}^{d}\left(\sum_{j=1}^na_r^{(j)}(a_s^{(j)})^*\right)\left(\sum_{k=1}^na_r^{(k)}(a_s^{(k)})^*\right)^*\\
	&\geq \sum_{r=1}^{d}\left(\sum_{j=1}^na_r^{(j)}(a_r^{(j)})^*\right)\left(\sum_{k=1}^na_r^{(k)}(a_r^{(k)})^*\right)^*\\
	&\geq \frac{1}{d}\left(\sum_{r=1}^{d}\left(\sum_{j=1}^na_r^{(j)}(a_r^{(j)})^*\right)\cdot 1\right)\left(\sum_{s=1}^{d}1\cdot\left(\sum_{k=1}^na_s^{(k)}(a_s^{(k)})^*\right)^*\right)\\
	&=\frac{1}{d}\left(\sum_{j=1}^n\sum_{r=1}^{d}a_r^{(j)}(a_r^{(j)})^* \right)\left(\sum_{k=1}^n\sum_{s=1}^{d}a_s^{(k)}(a_s^{(k)})^*\right)\\
	&=\frac{1}{d}\left(\sum_{j=1}^{n}1\right)\left(\sum_{k=1}^{n}1\right)=\frac{n^2}{d}.
\end{align*}
Hence 
\begin{align*}
	(n-1) a+1\geq   \frac{n}{d} \implies 	d(n-1) a+d\geq n \implies d(1-a)\geq n(1-da).
\end{align*}
\end{proof}
\begin{theorem} 
Let $\mathcal{A}$ be a commutative unital C*-algebra.	Let $\{\tau_j\}_{j=1}^n$  	be modular  $\gamma$-norm-equiangular  lines in $\mathcal{A}^d$. Then 
		\begin{align*}
		n(1-d\gamma^2) \leq d(1-\gamma^2).
	\end{align*}
	In particular, if 
		\begin{align*}
		\gamma < \frac{1}{\sqrt{d}},
	\end{align*}
	then 
	\begin{align*}
		n\leq \frac{d(1-\gamma^2)}{1-d\gamma^2}.
	\end{align*}
\end{theorem}
\begin{proof}
	By using the proof of Theorem \ref{MVS}, 
	\begin{align*}
		 \frac{n^2}{d}&\leq \sum_{j=1}^n\sum_{k=1}^n\langle \tau_j, \tau_k\rangle \langle \tau_k, \tau_j\rangle = \sum_{j,k=1, j\neq k}^n\langle \tau_j, \tau_k\rangle \langle \tau_k, \tau_j\rangle +\sum_{j=1}^n\langle \tau_j, \tau_j\rangle^2\\
		 &=\sum_{j,k=1, j\neq k}^n\langle \tau_j, \tau_k\rangle \langle \tau_k, \tau_j\rangle +n\leq \sum_{j,k=1, j\neq k}^n\|\langle \tau_j, \tau_k\rangle \|^2 +n=(n^2-n)\gamma^2+n.
	\end{align*}
\end{proof}

Proof of previous theorems used commutativity of C*-algebra. We are unable to derive noncommutative versions of them. However,  we introduce a subclass of norm-equiangular lines for which we derive modular relative bound.
\begin{definition}
	Let $\gamma\in [0,1].$   A collection $\{\tau_j\}_{j=1}^n$ in  $\mathcal{A}^d$  is said to be \textbf{special modular $\gamma$-norm-equiangular  lines} if following conditions hold.
	\begin{enumerate}[\upshape(i)]
		\item $	\langle \tau_j, \tau_j \rangle =1,  \forall 1\leq j \leq n$.
		\item $	\|\langle\tau_j, \tau_k \rangle\|=\gamma,  \forall 	1\leq j , k \leq n , j\neq k.$
		\item 
		\begin{align*}
			n^2 \leq d \sum_{j=1}^n\sum_{k=1}^n\langle \tau_j, \tau_k\rangle \langle \tau_k, \tau_j\rangle.
		\end{align*}
	\end{enumerate}
\end{definition}
\begin{theorem} (\textbf{Noncommutative van Lint-Seidel Relative Bound})
Let $\mathcal{A}$ be a  unital C*-algebra.	Let $\{\tau_j\}_{j=1}^n$  	be special modular  $\gamma$-norm-equiangular  lines in $\mathcal{A}^d$. Then 
\begin{align*}
	n(1-d\gamma^2) \leq d(1-\gamma^2).
\end{align*}
In particular, if $\gamma<\sqrt{d^{-1}}$, then 
\begin{align*}
	n\leq \frac{d(1-\gamma^2)}{1-d\gamma^2}.
\end{align*}
\end{theorem}
\begin{proof}
\begin{align*}
	n^2 \leq d \sum_{j=1}^n\sum_{k=1}^n\langle \tau_j, \tau_k\rangle \langle \tau_k, \tau_j\rangle\leq (n^2-n)\gamma^2+n.
\end{align*}	
\end{proof}
Next we derive modular Gerzon bound for modules over C*-algebras with invariant basis number  (IBN) property. These C*-algebras are studied in \cite{GIPSON}.
\begin{theorem} (\textbf{Modular Gerzon Universal Bound}) 
Let $\mathcal{A}$ be a  unital C*-algebra with IBN. Let $\{\tau_j\}_{j=1}^n$  	be  modular $a$-equiangular  lines in $\mathcal{A}^d$. If $1-a$ is invertible, then
		\begin{align}\label{MG}
			n \leq d^2.
		\end{align}	
\end{theorem}
\begin{proof}
For $1\leq j \leq n$, define 
\begin{align*}
	\tau_j \otimes \tau_j: \mathcal{A}^d \ni x \mapsto(\tau_j \otimes \tau_j)x\coloneqq\langle x, \tau_j\rangle \tau_j \in \mathcal{A}^d.
\end{align*}
We show that the collection $\{\tau_j \otimes \tau_j\}_{j=1}^n$   is linearly independent  over  $\mathcal{A}$. Let $c_1, \dots,  c_n \in \mathcal{A}$ be such that 
\begin{align}\label{12}
	\sum_{j=1}^{n}c_j(\tau_j \otimes \tau_j)=0.
\end{align}
Let $1\leq k \leq n$ be fixed. Then previous equation gives 
\begin{align*}
		0=\sum_{j=1}^{n}c_j(\tau_j \otimes \tau_j)\tau_k=\sum_{j=1}^{n}c_j\langle \tau_k, \tau_j\rangle \tau_j.
\end{align*}
Taking inner product with $\tau_k$ gives 
\begin{align*}
	0&=\sum_{j=1}^{n}c_j\langle \tau_k, \tau_j\rangle \langle \tau_j, \tau_k \rangle=\sum_{j=1, j\neq k}^nc_j\langle \tau_k, \tau_j\rangle \langle \tau_j, \tau_k\rangle +c_k\langle \tau_k, \tau_k\rangle^2\\
	&=\sum_{j=1, j\neq k}^nc_ja+c_k=\sum_{j=1}^nc_ja-c_ka+c_k=\sum_{j=1}^nc_ja+c_k(1-a).
\end{align*}
Now define 
\begin{align*}
	c\coloneqq \frac{-1}{1-a}\sum_{j=1}^{n}c_ja.
\end{align*}
Then $c_k=c$ for all $1\leq k \leq n$. We wish to show that $c=0$. Now Equation (\ref{12}) gives 
\begin{align*}
	0=	\sum_{j=1}^{n}c(\tau_j \otimes \tau_j).
\end{align*}
Let $1\leq k \leq n$ be fixed. Then previous equation gives 
\begin{align*}
	0=\sum_{j=1}^{n}c(\tau_j \otimes \tau_j)\tau_k=\sum_{j=1}^{n}c\langle \tau_k, \tau_j\rangle \tau_j.
\end{align*}
Taking inner product with $\tau_k$ gives 
\begin{align*}
	0&=\sum_{j=1}^{n}c\langle \tau_k, \tau_j\rangle \langle \tau_j, \tau_k \rangle=\sum_{j=1, j\neq k}^nc\langle \tau_k, \tau_j\rangle \langle \tau_j, \tau_k\rangle +c\langle \tau_k, \tau_k\rangle^2\\
	&=\sum_{j=1, j\neq k}^nca+c=(n-1)ca+c=c((n-1)a+1).
\end{align*}
Since $a\geq0$, $((n-1)a+1)$ is invertible, hence $c=0$. Note that  the matrix of $\tau_j \otimes \tau_j$ is $\tau_j \tau_j^*$ (viewing $\tau_j$ as column vector). The rank of $d$ by $d$ matrices over $\mathcal{A}$ is $d^2$. Since $\{\tau_j \otimes \tau_j\}_{j=1}^n$ is linearly independent and $\mathcal{A}$ has IBN property, $n\leq d^2$. 
\end{proof}
We can slightly generalize Definition \ref{A}. In the vector space case, such a generalization was done by Greaves,  Iverson, Jasper and Mixon \cite{GREAVESIVERSONJASPERMIXON}.
\begin{definition}\label{AB}
Let $a, b \in \mathcal{A}$ be  positive elements.   A collection $\{\tau_j\}_{j=1}^n$ in  $\mathcal{A}^d$  is said to be \textbf{modular $(a,b)$-equiangular  lines} if following conditions hold.
\begin{enumerate}[\upshape(i)]
	\item $	\langle \tau_j, \tau_j \rangle =b,  \forall 1\leq j \leq n$.
	\item $	\langle\tau_j, \tau_k \rangle\langle\tau_k, \tau_j \rangle=a,  \forall 	1\leq j , k \leq n , j\neq k.$
\end{enumerate}	
\end{definition}
Note that $b$ is necessarily positive but need not be invertible. Thus we may not able to reduce Definition \ref{AB} to Definition \ref{A}. By modifying earlier proofs, we easily get following theorems. 
\begin{theorem}
Let $\mathcal{A}$ be a commutative unital C*-algebra. 	Let $\{\tau_j\}_{j=1}^n$  	be modular  $(a,b)$-equiangular  lines in $\mathcal{A}^d$.
Then 
\begin{align*}
	n(b^2-da) \leq d(b-a).
\end{align*}
In particular, if $b^2-da$ is invertible, then 
\begin{align*}
	n\leq \frac{d(b-a)}{b^2-da}\leq d \left\|\frac{b-a}{b^2-da}\right\|.
\end{align*}	
\end{theorem}
\begin{theorem}
	Let $\mathcal{A}$ be a  unital C*-algebra with IBN. Let $\{\tau_j\}_{j=1}^n$  	be  modular $(a,b)$-equiangular  lines in $\mathcal{A}^d$. If $b^2-a$ and $(n-1)a+b^2$ are invertible, then
	\begin{align*}
		n \leq d^2.
	\end{align*}	
\end{theorem}
Similar to scalar  equiangular lines and Zauner conjecture, we formulate following problem.
\begin{problem}\label{NZ}
Let $\mathcal{A}$ be a  unital C*-algebra.	For which parameters $d, n \in \mathbb{N}$ and $a \in \mathcal{A}$, there exists a collection $\{\tau_j\}_{j=1}^n$  of  modular $a$-equiangular lines in $\mathcal{A}^d$? In particular, for which $d$/for all $d$, there is $d^2$ modular $\frac{1}{d+1}$-equiangular lines in $\mathcal{A}^d$ (making equality in Inequality (\ref{MG}))?
\end{problem}
Note that Problem \ref{NZ} contains Zauner conjecture (whenever $\mathcal{A}=\mathbb{C}$).

\section{Acknowledgments}
This paper has been partially developed when the author attended the workshop “Quantum groups, tensor categories and quantum field theory”, held in the University of Oslo, Norway from January 13 to 17, 2025. This event was organized by the University of Oslo, Norway and funded by the Norwegian Research Council through the “Quantum Symmetry” project.

 \bibliographystyle{plain}
 \bibliography{reference.bib}

\end{document}